Sally is 4 years older than Tim. Together, the sum of their ages is 20. How old is each?

Answer by JimboP1977(311)

Let S = Sally's age, T = Tim's age
S = T+4
S+T = 20
Subsitute T+4 into the second equation:
T+4+T = 20
2T+4 = 20
2T = 16
T = 8
So Tim is 8 and Sally is 12.